Teradene Corporation purchased land as a factory site and contracted with Maxtor Construction to construct a factory. Teradene made the following expenditures related to the acquisition of the land, building, and equipment for the factory: (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.): Purchase price of the land $ 1,330,000 Demolition and removal of old building 93,000 Clearing and grading the land before construction 215,000 Various closing costs in connection with acquiring the land 55,000 Architect's fee for the plans for the new building 63,000 Payments to Maxtor for building construction 3,380,000 Equipment purchased 925,000 Freight charges on equipment 45,000 Trees, plants, and other landscaping 58,000 Installation of a sprinkler system for the landscaping 6,300 Cost to build special platforms and install wiring for the equipment 25,000 Cost of trial runs to ensure proper installation of the equipment 8,300 Fire and theft insurance on the factory for the first year of use 37,000 In addition to the above expenditures, Teradene purchased four forklifts from Caterpillar. In payment, Teradene paid $29,000 cash and signed a noninterest-bearing note requiring the payment of $83,000 in one year. An interest rate of 6% properly reflects the time value of money for this type of loan. Required: Determine the initial valuation of each of the assets Teradene acquired in the above transactions. (Round your answers to the nearest whole dollars.)
Answer #1

Initial valuation of land = 1,330,000+93000+215000+55000 = 1,693,000

Initial valuation of building = 63000+3,380,000 = 3,443,000

Initial valuation of equipment = 925000 + 45000 + 25000 + 8300 = 1,003,300

Initial valuation of land improvement = 58000 + 6300 = 64300

Prepaid insurance = 37000

Initial valuation of forklift = 29000 + 83000 - [ 83000*6/106] =107302